An executive secretary can finish a typing job in 10 hours while an ordinary employee can do the same job in 12 hours. How long will it take the two employees to finish the typing if they work together?
Answer by richwmiller(17219)

1/10+ 1/12=1/x
6/60+5/60=1/x
11/60=1/x
x=60/11